One mole of N2 and 3 moles of PCl5 are placed in a 100 litre vessel heated to 227ºC. The equilibrium pressure is 2.05 atm. Assuming ideal behaviour, calculate degree of dissociation of PCl5 and KP 

Grade:12

2 Answers

Vikas TU
7 years ago
Write the eqn. first,
PCl5 ------->   PCl3   +  Cl2
Use ideal gas eqn. to get the value of n .i.e moles 
and use concenrtrn for findoong Kpinstead of using partial pressure.
Kp  = pcl3*cl2/pcl5
jatin meena
5 years ago
                 PCL5 PCL3 + CL2                                                                                                                                   3 mol          0 mol      0 mol  no.moles initially                                                                                                             3(3-x)            3x          3x  no.of moles at equi.                                                                                                   now,     
     on using 
                                 PV=nRT [P=2.05 , V=100 , R=0.0821 , T=227 DEGREE CECIOUS MEANS 500K]
                    =>   2.05*100=n*0.0821*500
                    =>  205=41.05n 
                      =>  n=4.9[ we can take approx 5]
                     total number of moles at equilirium 
           =>3-3x+3x+3x+1=5[we have taken 1  here because we have to add no. of moles of N2]
   => 3x+4=5
       => x=0.33[degree of dissociation] 
frome here we can find Kp also that will be 0.205atm